A wheel whose radius measures 16 inches is rotated. If a point on the circumference of the wheel moves through an arc of 12 feet what is the measure in radians of the angle through which a spoke of the wheel travels
Answer by TimothyLamb(4379)

c = 2*pi*r
c = 2*3.1415*(16/12)
c = 8.37733333333 feet
---
arc length in circumferences:
12/c = 12/8.37733333333 = 1.43243673404 circumferences
---
arc length in radians:
2pi radians/c
---
1.43243673404 c * (2*3.1415 radians/c) = 9 radians
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---